OpenStudy (anonymous):

Four times a number plus 8 results in 15. What is the number? Answer x = 7/4 x = 19 x = 23/4 x = 27

OpenStudy (anonymous):

@ktnguyen1

OpenStudy (anonymous):

7/4

OpenStudy (anonymous):

4n+8= 15 4n=15-8 n=7divided by 4